find the general equation of a line parallel to the line 2x+3y-6=0 and passing at a distance 5 square root 13 over 13 from the point (-1,1)
1
Expert's answer
2020-12-01T06:30:03-0500

Let us find the general equation of a line parallel to the line "2x+3y-6=0" and passing at a distance "5\frac{\sqrt{13}}{13}" from the point "(-1,1)".


The vector "\overrightarrow{n}=(2,3)" is perpendicular to all lines parallel to "2x+3y-6=0" . Thererfore, the euqation of any line parralel to "2x+3y-6=0" is "2x+3y+c=0" for some "c\in\mathbb R".


The distance from the point "(-1,1)" to the line "2x+3y+c=0" is "d=\frac{|2(-1)+3\cdot 1+c|}{\sqrt{4+9}}=\frac{|1+c|}{\sqrt{13}}=|1+c|\frac{\sqrt{13}}{13}". Onthe other hand "d=5\frac{\sqrt{13}}{13}". Therefore, "|1+c|=5", and hence "1+c=\pm 5". So, "c=-6" or "c=4." We conclude that there are two lines parallel to the line "2x+3y-6=0" and passing at a distance "5\frac{\sqrt{13}}{13}" from the point "(-1,1)": "2x+3y-6=0" and "2x+3y+4=0". One of them coinside with the initial line "2x+3y-6=0" .